# Benchmarking Performance, Explainability, and Evaluation Strategies of Vision-Language Models for Surgery: Challenges and Opportunities

### v1

@@ 5.3 Discussion
Based on the results, we observe that the overall performance of vision-language models (VLMs) on surgical data is generally limited. For CLIP, we observed that it can focus on the correct object and achieves relatively good ARS scores for the verb associated with the instrument. This suggests that vision-language models (VLMs) trained on diverse, large-scale datasets can still perform relatively well on downstream tasks, leveraging their generalized understanding from broad visual concepts. However, we also observed a notable gap between classification accuracy and ARS scores, indicating that the models often make correct predictions without actually attending to the correct spatial regions.
@@ 6 Conclusion
Our work introduces a benchmarking framework for evaluating vision-language models (VLMs) in surgical settings, with a focus on spatial attention and grounding beyond standard accuracy metrics. We propose new evaluation measures that assess how well model attention aligns with corresponding relevant regions, providing deeper insight into model behavior. Our findings reveal not only the generally low performance of VLMs on surgical data, but also a consistent disconnect between correct predictions and appropriate visual focus. These issues raise critical concerns about the reliability and interpretability of current VLMs in surgical applications. Even models trained specifically on surgical data struggle when faced with frame-level, fine-grained tasks such as action recognition and instrument classification.

### v2

@@ 5.3 Discussion
Based on the results, we observe that the overall performance of VLMs on surgical data is generally limited. For CLIP, we observed that it can focus on the correct object and achieves relatively good ARS scores for the verb associated with the instrument. This suggests that VLMtrained on diverse, large-scale datasets can still perform relatively well on downstream tasks, leveraging their generalized understanding from broad visual concepts. However, we also observed a notable gap between classification accuracy and ARS scores, indicating that the models often make correct predictions without actually attending to the correct spatial regions.
@@ 6 Conclusion
Our work introduces a benchmarking framework for evaluating VLMs in surgical settings, with a focus on spatial attention and grounding beyond standard accuracy metrics. We propose new evaluation measures that assess how well model attention aligns with corresponding relevant regions, providing deeper insight into model behavior. Our findings reveal not only the generally low performance of VLMs on surgical data, but also a consistent disconnect between correct predictions and appropriate visual focus. These issues raise critical concerns about the reliability and interpretability of current VLMs in surgical applications. Even models trained specifically on surgical data struggle when faced with frame-level, fine-grained tasks such as action recognition and instrument classification.